Competition

Assume the following:
Qd = 620 – 5p
Qs = 170 + 5p
TC= 5 + 4Q + 2Q2
Find P, Q that max profit then TR, TC, and profit.
Answer #1

Answer

the market equilibrium is at Qd=Qs

620-5P=170+5P

10P=450

P=45

Q=620-5*45

=395

the market quantity is 395

===

the firm produces at MC=P

MC=change in total cost =first differentiation of total cost =dTC/dQ=4+4Q

4+4Q=45

4Q=41

Q=10.25 ......... the quantity maximize profit is 10.25, and P=45

TR=P*Q=10.25*45=461.25

TC=5+4*10.25+2*10.25^2

=256.125

Profit=TR-TC

=461.25-256.125

=205.125

profit is $205.125